OpenStudy (anonymous):

PLEASE HELP!!! In ELECTRONICS, the total resistance to the flow of electricity in a circuit is called the impedance, Z. Impedance is represented by a complex number. The total impedance in a series circuit is the sum of individual imepedances. The impedance in one part of a circuit is Z1 = 3 + 4i. In another part of a circuit, the impedance is Z1 = 5 - 2i. What is the total impedance of the circuit?

OpenStudy (anonymous):

sorry it call Imaginary Numbers

OpenStudy (anonymous):

8+2i

OpenStudy (anonymous):

Assuming the impedances are connected in series, total impedance is Z=Z1+Z2 Z=(3+4i)+(5-2i)=(3+5)+(4i-2i)=8+2i

OpenStudy (anonymous):

Thank u!!!